OpenStudy (anonymous):

Write the trigonometric function values in terms of its cofunction sin(60degrees-x)

OpenStudy (anonymous):

cos [90-(60-x)] = cos(90-60+x) = cos (30+x)

OpenStudy (anonymous):

Hello could you please tell me how does it become Cosine? and with equation like that?

OpenStudy (anonymous):

ok sin(90-A) = cosA cos(90-A) = sinA